# Further study of $c\bar{c}c\bar{c}$ system within a chiral quark model

### 15 Mar 2024 v1

@@ Abstract
Inspired by the recent Altas and CMS experiments on the invariant mass spectrum of $J/\psi J/\psi$, we systematically study the $c\bar{c}c\bar{c}$ system of $J^{P}=0^{+}$. In the framework of chiral quark model, we have carried out bound-state calculation and resonance-state calculation respectively by using Real-scaling method. The results of bound-state calculation show that there are no bound states in the $c\bar{c}c\bar{c}$ with $0^{+}$ system. The resonance-state calculation shows that there are four possible stable resonances: $R(6920)$, $R(7000)$, $R(7080)$ and $R(7160)$. $R(6920)$ and $R(7160)$ are experimental candidates for $X(6900)$ and $X(7200)$, whose main decay channel is $J/\psi J/\psi$. It is important to note that the another major decay channel of $R(7160)$ is $\chi_{c0}\chi_{c0}$, and the $\chi_{c0}\chi_{c0}$ is also the main decay channel of $R(7000)$, $R(7080)$. Therefore, we propose to search experimentally for these two predicted resonances in the $\chi_{c0}\chi_{c0}$ invariant mass spectrum.
@@ I Introduction
Subsequently, CMS CMS:2023owd and ATLAS ATLAS:2023bft Collaborations confirmed the existence of $X(6900)$. In addition, some new resonances (R) are observed by CMS and ATALS Collaborations. The mass and decay width of these resonance structure from CMS are
@@
While for ATLAS, in the di-$J/\psi$ and $J/\psi$$\psi(2S)$ channels, several resonances are
@@
After the experimental reports, the study of fully-charm tetraquark states arose again. In Ref. Jin:2020jfc , Jin et al. employed two models, quark delocalization color screening model and chiral quark model, to study the fully-charm tetraquark system, and the results are qualitative consistent in two quark models. The $X(6900)$ state was explained as a $S$-wave compact resonance state with $J^{P}=0^{+}$. Besides, the authors also found resonance structures around the energies $6270$ MeV with $J^{P}=0^{+}$ and $7260$ MeV with $J^{P}=0^{+}$ and $1^{+}$, respectively. However, no signal was found for $X(6600)$ state. Li et al. Li:2021ygk applied the Bethe-Salpeter equation to investigate the states of fully charm tetraquark, and they obtained three ground states of $cc\bar{c}\bar{c}$ in the mass range from $6.4$ GeV to $6.5$ GeV. For $X(6900)$, they suggested that it is less likely to be the ground state of a compact $cc\bar{c}\bar{c}$ tetraquark, but may be the first or second excited states. In Ref. Wang:2022yes , Wang et al. calculated the $S$-wave $cc\bar{c}\bar{c}$ tetraquark states in the framework of nonrelativistic quark model. Their results show that in the $S$-wave sector, no resonance is found at the energy region of the $X(6200)$ and $X(6600)$ states. Although for $X(6900)$ they obtained masses with $J^{PC}=0^{++}$ and $2^{++}$ obove the experimental $100$ MeV, the decay widths are in agreement with the experiment. Moreover, the higher $0^{++}$ and $2^{++}$ resonances are found at about $7.2$ GeV with the widths of $60.6$ MeV and $91.2$ MeV. More results and discussions on $S$-wave for $cc\bar{c}\bar{c}$ tetraquark systems are given in Refs. Yang:2021hrb ; Lu:2020cns ; Yang:2020wkh ; Liang:2021fzr ; Dong:2020nwy ; Wan:2020fsk ; Feng:2020riv ; Maciula:2020wri ; Feng:2020qee ; Goncalves:2021ytq ; Huang:2021vtb ; Becchi:2020uvq ; Bedolla:2019zwg ; Faustov:2020qfm ; Gong:2020bmg ; Gordillo:2020sgc ; Karliner:2020dta ; Ke:2021iyh ; Weng:2020jao ; Zhang:2020xtb ; Zhao:2020jvl ; Zhao:2020nwy ; Wu:2022qwd ; Zhuang:2021pci ; Wang:2022xja .
For the fully-heavy tetraquark states, the theoretical explanations are far from convergent. By using QCD sum rule, Ref. Chen:2020xwe argued that the $X(6900)$ can be considered as a $P$-wave $cc\bar{c}\bar{c}$ tetraquark with $J^{PC}=0^{-+}$ or $1^{-+}$. In Ref. Zhu:2020xni , in the framework of Bethe-Salpeter and Regge trajectory relation, the author pointed out that the $X(6900)$ can be explained as a radially excited state with $J^{PC}=0^{++}(3S)$ or $J^{PC}=2^{++}(3S)$ or an orbitally excited $2P$ state. Deng et al. Deng:2020iqw pointed out that the structure $X(6900)$ can be identified as the excited $[cc][\bar{c}\bar{c}]$ with $L=1~{}(L=2)$ in the color flux-tube model. In Ref. Zhang:2021sio , the authors suggested that $X(6900)$ may be caused by $1P$-, or $2S$-, or $1D$-wave states while the $X(7200)$ is the high-lying $1D$-wave states with $J^{PC}=0^{++},1^{++},2^{++},3^{++}$, or $4^{++}$.

### 21 Mar 2024 v2

@@ Abstract
Inspired by the recent ATLAS and CMS experiments on the invariant mass spectrum of $J/\psi J/\psi$, we systematically study the $c\bar{c}c\bar{c}$ system of $J^{P}=0^{+}$. In the framework of chiral quark model, we have carried out bound-state calculation and resonance-state calculation respectively by using Real-scaling method. The results of bound-state calculation show that there are no bound states in the $c\bar{c}c\bar{c}$ with $0^{+}$ system. The resonance-state calculation shows that there are four possible stable resonances: $R(6920)$, $R(7000)$, $R(7080)$ and $R(7160)$. $R(6920)$ and $R(7160)$ are experimental candidates for $X(6900)$ and $X(7200)$, whose main decay channel is $J/\psi J/\psi$. It is important to note that the another major decay channel of $R(7160)$ is $\chi_{c0}\chi_{c0}$, and the $\chi_{c0}\chi_{c0}$ is also the main decay channel of $R(7000)$, $R(7080)$. Therefore, we propose to search experimentally for these two predicted resonances in the $\chi_{c0}\chi_{c0}$ invariant mass spectrum.
@@ I Introduction
Subsequently, CMS CMS:2023owd and ATLAS ATLAS:2023bft Collaborations confirmed the existence of $X(6900)$. In addition, some new resonances (R) are observed by CMS and ATALS Collaborations. The mass and decay width of these resonance structures from CMS are:
@@
While for ATLAS, in the di-$J/\psi$ and $J/\psi$$\psi(2S)$ channels, several resonances are:
@@
After the experimental reports, the study of fully-charm tetraquark states arose again. In Ref. Jin:2020jfc , Jin et al. employed two models, quark delocalization color screening model and chiral quark model, to study the fully-charm tetraquark system, and the results are qualitative consistent in two quark models. The $X(6900)$ state was explained as a $S$-wave compact resonance state with $J^{P}=0^{+}$. Besides, the authors also found resonance structures around the energies $6270$ MeV with $J^{P}=0^{+}$ and $7260$ MeV with $J^{P}=0^{+}$ and $1^{+}$, respectively. However, no signal was found for $X(6600)$ state. Li et al. Li:2021ygk applied the Bethe-Salpeter equation to investigate the states of fully charm tetraquark, and they obtained three ground states of $cc\bar{c}\bar{c}$ in the mass range from $6.4$ GeV to $6.5$ GeV. For $X(6900)$, they suggested that it is less likely to be the ground state of a compact $cc\bar{c}\bar{c}$ tetraquark, but may be the first or second excited states. In Ref. Wang:2022yes , Wang et al. calculated the $S$-wave $cc\bar{c}\bar{c}$ tetraquark states in the framework of nonrelativistic quark model. Their results show that in the $S$-wave sector, no resonance is found at the energy region of the $X(6200)$ and $X(6600)$ states. Although for $X(6900)$ they obtained masses with $J^{PC}=0^{++}$ and $2^{++}$ obove the experimental $100$ MeV, the decay widths are in agreement with the experiment. Moreover, the higher $0^{++}$ and $2^{++}$ resonances are found at about $7.2$ GeV with the widths of $60.6$ MeV and $91.2$ MeV. In Ref. Wu:2024euj , Wu et al. investigated the $S$-wave fully heavy tetraquark systems by utilizing the Gaussian expansion method. They found $X(6900)$ and $X(7200)$ can be explained as $0^{++}$ and $2^{++}$ systems. However, $X(6400)$ and $X(6600)$ were not seen in their calculation.
[…]